please help me with these math problems
6z / 5x = 30
10x+20y=40
4x+7y=28

Answer by Alan3354(69443)

please help me with these math problems
6z / 5x = 30
10x+20y=40
4x+7y=28
===============
1 step at a time.
Step 1, what do you want to do?